OpenStudy (anonymous):

Find and simplify f(-x)-f(x) for the given function f. f(x)=-7x^3-3x^2+9x-20. f(-x)-f(x)=?

jimthompson5910 (jim_thompson5910):

f(x)=-7x^3-3x^2+9x-20 f(-x)=-7(-x)^3-3(-x)^2+9(-x)-20 f(-x)=-7(-x^3)-3(x^2)+9(-x)-20 f(-x)=7x^3-3x^2-9x-20 -------------------------------------- f(-x)-f(x) (7x^3-3x^2-9x-20) - (-7x^3-3x^2+9x-20) 7x^3-3x^2-9x-20 + 7x^3+3x^2-9x+20 14x^3-18x So f(-x)-f(x) = 14x^3-18x